A Bibliography of Classic Shell Scripting

Arnold Robbins

and

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

FAX: +1 801 585 1640, +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

26 November 2011 Version 1.04

Title word cross-reference

1003.1-2001 [19, 20, 21, 22].

4.4BSD [35].

6 [19, 20, 21]. 6th [34].

7 [44].

Accidental [41, 42]. Advanced [55, 62]. Algorithms [29, 51]. Applied [51]. Architecture [18]. Art [43]. awk [46, 1, 45, 11].

Base [19]. bash [38, 39]. Bazaar [41, 42]. Beyond [60]. book [54].

C [51]. Cathedral [41, 42]. century [50]. Classic [47]. Co [60]. Code [34, 51, 54]. Coder [6]. Command [9]. Commentary [34]. Commerce [16]. Common [56]. Confessions [6]. corrections [7]. CRT [32]. Cryptographic [18]. Cryptography [51, 54]. csh [12]. CUPS [56].

Definitions [19]. Definitive [4, 37]. Design [3, 18, 35, 36]. Desktop [44, 28, 53]. Digital [52]. Displays [32]. Documentation [10]. Domain [32].

Eavesdropping [32]. Edition [34]. Editor [33]. Effective [45]. Efficient [8]. Elements [24]. Environment [27, 55, 62]. evolution [54]. Expressions [13].

Father [60]. Forgotten [2]. Format [10]. FreeBSD [36].

Giant [2]. GNU [10]. Good [15]. Guide [4, 37, 61].

Hands [48]. Hands-On [48]. Herman [2]. Hollerith [2]. HTML [37].

IBM [60]. Implementation [35, 36]. Information [2, 19, 20, 21, 22]. Informative [22]. Interface [19, 20, 21, 22]. Interfaces [20]. Internals [58]. Internet [17]. Issue [19, 20, 21].

Korn [49]. KornShell [9]. KornShell93 [48].

Language [1, 9]. Learning [33, 38, 39, 40, 49]. Lies [52]. Life [60]. Linux [14, 41, 42, 53]. Lions’ [34]. Literate [30].

Making [59]. Mary [54]. Mastering [13]. Musings [41, 42]. My [60].

Networked [52]. Nutshell [44, 28, 53].

Official [61]. open [42, 41]. Operating [3, 19, 20, 21, 22, 35, 36, 40]. Optical [32].

Pascal [26]. Pearls [6, 7, 5]. PGP [15, 61]. Philosophy [14]. Pocket [46]. Portable [19, 20, 21, 22]. POSIX [19, 20, 21, 22]. Practical [17]. Practice [23]. Pretty [15]. Printing [56]. Privacy [15, 16]. Processing [2]. Programming [1, 6, 7, 5, 9, 24, 27, 23, 30, 43, 45, 48, 55, 62]. Programs [8]. Protocols [51].

quantum [54]. quarter [50]. Queen [54]. Quick [44, 28, 53].

Rationale [22]. Reference [44, 46, 28, 53]. Regular [13]. reprinted [7]. Revolutionary [41, 42]. Risks [32].

Scots [54]. Scripting [47]. secrecy [54]. Secrets [52]. Secure [4]. Security [16, 17, 18, 52]. sed [46, 11]. Seminumerical [29]. September [10]. Shell [21, 38, 47, 49, 39, 4]. Software [25, 26]. Solaris [44]. Son [60]. Source [34, 51, 42, 41]. SQL [28]. SSH [4]. Standard [19, 20, 21, 22, 57]. Std [19, 20, 21, 22]. Style [24]. SVR4 [44]. System [3, 19, 20, 21, 22, 35, 36, 40, 56].

tcsh [12]. Technology [19, 20, 21, 22]. TEXbook [31]. TeXinfo [10]. Time [32]. Time-Domain [32]. Tools [25, 26].

Unicode [57]. UNIX [56, 3, 27, 34, 40, 43, 44, 50, 55, 58, 14, 17, 62]. User [61]. Using [12]. Utilities [21].

Verification [18]. Version [57, 10]. vi [33].

Web [16]. Work [59]. World [52]. Writing [8].

XHTML [37].

References

Aho:1988:APL

[1] Alfred V. Aho, Brian W. Kernighan, and Peter J. Weinberger. The AWK Programming Language. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-07981-X. x + 210 pp. LCCN QA76.73.A95 A35 1988.

Austrian:1982:HHF

[2] Geoffrey D. Austrian. Herman Hollerith — Forgotten Giant of Information Processing. Columbia University Press, New York, NY, USA, 1982. ISBN 0-231-05146-8. xvi + 418 pp. LCCN QA76.2.H64 .A97.

Bach:1986:DUO

[3] Maurice J. Bach. The Design of the UNIX Operating System. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1986. ISBN 0-13-201799-7. xiv + 471 pp. LCCN QA76.76.O63 B33 1986.

Barrett:2001:SSS

[4] Daniel J. Barrett and Richard E. Silverman. SSH: The Secure Shell: The Definitive Guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2001. ISBN 0-596-00011-1. xv + 540 pp. LCCN QA76.76.O63 B369 2001. US$39.95. URL http://www.snailbook.com/;

     http://www.oreilly.com/catalog/sshtdg/.

Bentley:2000:PP

[5] Jon Louis Bentley. Programming Pearls. Addison-Wesley, Reading, MA, USA, second edition, 2000. ISBN 0-201-65788-0. xi + 239 pp. LCCN QA76.6 .B454 2000. US$24.95.

Bentley:1988:MPP

[6] Jon Louis Bentley. More Programming Pearls: Confessions of a Coder. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-11889-0. viii + 207 pp. LCCN QA76.6 .B452 1988. US$18.75.

Bentley:1989:PPR

[7] Jon Louis Bentley. Programming Pearls (reprinted with corrections). Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-10331-1. viii + 195 pp. LCCN QA76.6.B453 1989.

Bentley:1982:WEP

[8] Jon Louis Bentley. Writing Efficient Programs. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1982. ISBN 0-13-970251-2 (hardcover), 0-13-970244-X (paperback). xvi + 170 pp. LCCN QA76.6 .B455 1982. US$22.95 (hardcover), US$14.95 (paperback).

Bolsky:1995:NKC

[9] Morris I. Bolsky and David G. Korn. The New KornShell Command And Programming Language. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, second edition, 1995. ISBN 0-13-182700-6. 416 pp. LCCN QA76.73.K67 B64 1995. ????

Chassell:1999:TGD

[10] Robert J. Chassell and Richard M. Stallman. TeXinfo: The GNU Documentation Format (for Texinfo version 4.0, 28 September 1999). Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 1999. ISBN 1-882114-67-1. x + 244 pp. LCCN ???? US$25.00.

Dougherty:1997:SA

[11] Dale Dougherty and Arnold Robbins. sed and awk. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, February 1997. ISBN 1-56592-225-5. xxii + 407 pp. LCCN QA76.76.U84 D69

1997. US$29.95. URL http://www.oreilly.com/catalog/sed2.

DuBois:1995:UCT

[12] Paul DuBois. Using csh and tcsh. A Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, August 1995. ISBN 1-56592-132-1. xviii + 221 pp. LCCN QA76.76.O63 D82 1995.

US$29.95. URL http://www.oreilly.com/catalog/tcsh.

Friedl:2002:MRE

[13] Jeffrey E. F. Friedl. Mastering Regular Expressions. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2002. ISBN 0-596-00289-0. xxii + 460 pp. LCCN QA76.73.P22 F75 2002; QA76.9.T48 F75 2002. US$39.95. URL http://www.oreilly.com/

catalog/regex2.

Gancarz:2003:LUP

[14] Mike Gancarz. Linux and the Unix Philosophy. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 2003. ISBN 1-55558-273-7. xxvii + 220 pp. LCCN QA76.76.O63G364 2003. US$34.99. URL

http://www.loc.gov/catdir/description/els031/2003051482.html;
 http://www.loc.gov/catdir/toc/els031/2003051482.html.

Garfinkel:1995:PPG

[15] Simson Garfinkel. PGP: Pretty Good Privacy. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1995. ISBN 1-56592-098-8. xxxiii + 393 pp. LCCN QA76.9.A25G36 1995. US$24.95.

Garfinkel:2002:WSP

[16] Simson Garfinkel and Gene Spafford. Web Security, Privacy & Commerce. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2002. ISBN 0-596-00045-6. xxviii + 756 pp. LCCN TK5105.59 .G37 2002. US$44.95. URL http://safari.oreilly.

com/0596000456;   http://www.oreilly.com/catalog/websec2.

Garfinkel:2003:PUI

[17] Simson Garfinkel, Gene Spafford, and Alan Schwartz. Practical Unix & Internet Security. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2003. ISBN 0-596-00323-4. xxix + 954 pp. LCCN QA76.76.O63 G38 2003. US$54.95. URL http://www.oreilly.

com/catalog/puis3.

Gutmann:2004:CSA

[18] Peter Gutmann. Cryptographic Security Architecture: Design and Verification. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 0-387-95387-6. xviii + 320 pp. LCCN QA76.9.A25 G88 2002.

IEEE:2001:ISRa

[19] IEEE. IEEE Std 1003.1-2001 Standard for Information Technology — Portable Operating System Interface (POSIX) Base Definitions, Issue 6. IEEE, New York, NY, USA, 2001. ISBN 1-85912-247-7 (UK), 1-931624-07-0 (US), 0-7381-3047-8 (print), 0-7381-3010-9 (PDF), 0-7381-3129-6 (CD-ROM). xliv + 448 pp. LCCN ???? Revision of IEEE Std 1003.1-1996 and IEEE Std 1003.2-1992) Open Group Technical Standard Base Specifications, Issue 6.

IEEE:2001:ISRb

[20] IEEE. IEEE Std 1003.1-2001 Standard for Information Technology — Portable Operating System Interface (POSIX) System Interfaces, Issue 6. IEEE, New York, NY, USA, 2001. ISBN 1-85912-247-7 (UK), 1-931624-07-0 (US), 0-7381-3094-4 (invalid checksum??) (print), 0-7381-3010-9 (PDF), 0-7381-3129-6 (CD-ROM). xxx + 1690 pp. LCCN ???? Revision of IEEE Std 1003.1-1996 and IEEE Std 1003.2-1992, Open Group Technical Standard Base Specifications, Issue 6.

IEEE:2001:ISRc

[21] IEEE. IEEE Std 1003.1-2001 Standard for Information Technology — Portable Operating System Interface (POSIX) Shell and Utilities, Issue 6. IEEE, New York, NY, USA, 2001. ISBN 1-85912-247-7 (UK), 1-931624-07-0 (US), 0-7381-3050-8 (print), 0-7381-3010-9 (PDF), 0-7381-3129-6 (CD-ROM). xxxii + 1090 pp. LCCN ???? Revision of IEEE Std 1003.1-1996 and IEEE Std 1003.2-1992) Open Group Technical Standard Base Specifications, Issue 6.

IEEE:2001:ISRd

[22] IEEE. IEEE Std 1003.1-2001 Standard for Information Technology — Portable Operating System Interface (POSIX) Rationale (Informative). IEEE, New York, NY, USA, 2001. ISBN 1-85912-247-7 (UK), 1-931624-07-0 (US), 0-7381-3048-6 (print), 0-7381-3010-9 (PDF), 0-7381-3129-6 (CD-ROM). xxxiv + 310 pp. LCCN ???? Revision of IEEE Std 1003.1-1996 and IEEE Std 1003.2-1992) Open Group Technical Standard Base Specifications, Issue 6.

Kernighan:1999:PP

[23] Brian W. Kernighan and Rob Pike. The Practice of Programming. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-61586-X. xii + 267 pp. LCCN QA76.6 .K48 1999. US$24.95, CAN$37.50. URL http://tpop.awl.

com;   http://cseng.aw.com/bookdetail.qry?ISBN=0-201-61586-X&
ptype=0;   http://cm.bell-labs.com/cm/cs/tpop/code.html.

Kernighan:1978:EPS

[24] Brian W. Kernighan and P. J. Plauger. The Elements of Programming Style. McGraw-Hill, New York, NY, USA, second edition, 1978. ISBN 0-07-034207-5. xii + 168 pp. LCCN QA76.6 .K39 1978.

Kernighan:1976:ST

[25] Brian W. Kernighan and P. J. Plauger. Software Tools. Addison-Wesley, Reading, MA, USA, 1976. ISBN 0-201-03669-X. 338 pp. LCCN QA76.6 .K42 1976.

Kernighan:1981:STP

[26] Brian W. Kernighan and P. J. Plauger. Software Tools in Pascal. Addison-Wesley, Reading, MA, USA, 1981. ISBN 0-201-10342-7. ix + 366 pp. LCCN QA76.6 .K493. US$13.95.

Kernighan:1984:UPE

[27] Brian W. Kernighan and Rob Pike. The UNIX Programming Environment. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1984. ISBN 0-13-937699-2 (hardcover), 0-13-937681-X (paperback). x + 357 pp. LCCN QA76.76.O63 K48 1984.

Kline:2001:SND

[28] Kevin E. Kline and Daniel Kline. SQL in a nutshell: a desktop quick reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2001. ISBN 1-56592-744-3. viii + 214 pp. LCCN QA76.73.S67 K55 2001. US$29.95.

Knuth:1997:SA

[29] Donald E. Knuth. Seminumerical Algorithms, volume 2 of The Art of Computer Programming. Addison-Wesley, Reading, MA, USA, third edition, 1997. ISBN 0-201-89684-2. xiii + 762 pp. LCCN QA76.6 .K64 1997. US$52.75.

Knuth:1992:LP

[30] Donald E. Knuth. Literate Programming. CSLI Lecture Notes Number 27. Stanford University Center for the Study of Language and Information, Stanford, CA, USA, 1992. ISBN 0-937073-80-6 (paper), 0-937073-81-4 (cloth). xv + 368 pp. LCCN QA76.6.K644. US$24.95.

Knuth:1984:T

[31] Donald E. Knuth. The TEXbook. Addison-Wesley, Reading, MA, USA, 1984. ISBN 0-201-13448-9. ix + 483 pp. LCCN Z253.4.T47 K58 1984.

Kuhn:2002:OTD

[32] Markus Kuhn. Optical time-domain eavesdropping risks of CRT displays. In IEEE, editor, Proceedings: 2002 IEEE Symposium on Security and Privacy, 12–15 May, 2002, Berkeley, California, pages 3–18. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. ISBN 0-7695-1543-6. LCCN QA76.9.A25 I34 2002. URL http://

www.cl.cam.ac.uk/~mgk25/ieee02-optical.pdf. IEEE Computer Soci-

ety Order Number PR01543.

Lamb:1998:LVE

[33] Linda Lamb and Arnold Robbins. Learning the vi Editor. A Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, sixth edition, 1998. ISBN 1-56592-426-6. xvii + 327 pp. LCCN QA76.76.O63 L355 1998. US$24.95. URL http://www.oreilly.

com/catalog/vi6.

Lions:1996:LCU

[34] John Lions. Lions’ Commentary on UNIX 6th Edition, with Source Code. Computer classics revisited. Peer-to-Peer Communications, San Jose, CA 95164-0218, USA, 1996. ISBN 1-57398-013-7. 254 pp. LCCN ????

US$29.96; CAN$41.95. URL http://www.peer-to-peer.com/catalog/

opsrc/lions.html. With forewords by Dennis M. Ritchie and Ken Thompson. Prefatory notes by Peter H. Salus and Michael Tilson; a Historical Note by Peter H. Salus; and Appreciations by Greg Rose, Mike O’Dell, Berny Goodheart, Peter Collinson, and Peter Reintjes. Originally circulated as two restricted-release volumes: “UNIX Operating System Source Code Level Six”, and “A Commentary on the UNIX Operating System”.

McKusick:1996:DIO

[35] Marshall Kirk McKusick, Keith Bostic, Michael J. Karels, and John S. Quarterman. The Design and Implementation of the 4.4BSD Operating System. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-54979-4. xxvi + 580 pp. LCCN QA76.76.O63D4743 1996. US$44.95.

McKusick:2005:DIF

[36] Marshall Kirk McKusick and George V. Neville-Neil. The Design and Implementation of the FreeBSD Operating System. Addison-Wesley, Reading, MA, USA, 2005. ISBN 0-201-70245-2. xxviii + 683 pp. LCCN QA76.76.O63

M398745 2005. URL http://www.mckusick.com/FreeBSDbook.html.

Musciano:2000:HXD

[37] Chuck Musciano and Bill Kennedy. HTML and XHTML: The Definitive Guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, fourth edition, 2000. ISBN 0-596-00026-X. xvi + 655 pp. LCCN QA76.76.H94 M875 2000 Bar. US$34.95. URL http://www.

oreilly.com/catalog/html4.

Newham:1998:LBS

[38] Cameron Newham and Bill Rosenblatt. Learning the bash Shell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 1998. ISBN 1-56592-347-2. xvi + 318 pp. LCCN QA76.76.O63

N458 1998. US$29.95. URL http://www.oreilly.com/catalog/bash2.

Newham:2005:LBS

[39] Cameron Newham and Bill Rosenblatt. Learning the bash shell. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, third edition, 2005. ISBN 0-596-00965-8 (paperback). xvi + 333 pp. £24.95.

Peek:2001:LUO

[40] Jerry D. Peek, Grace Todino, and John Strang. Learning the UNIX operating system. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, fifth edition, 2001. ISBN 0-596-00261-0. xiii + 157 pp. LCCN QA76.76.O63 T62 2002.

Raymond:1999:CBP

[41] Eric S. Raymond. The Cathedral and the Bazaar: Musings on Linux and Open Source by an Accidental Revolutionary. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, February 1999. ISBN 0-596-00108-8. 208 pp. LCCN QA76.76.O63 R396 1999. US$16.95. URL

http://www.oreilly.com/catalog/cathbazpaper.

Raymond:2001:CBM

[42] Eric S. Raymond. The cathedral and the bazaar: musings on Linux and open source by an accidental revolutionary. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, revised edition, 2001. ISBN 0-596-00131-2. xiv + 241 pp. LCCN QA76.76.O63 R397 2001.

Raymond:2004:AUP

[43] Eric Steven Raymond. The Art of UNIX Programming. Addison-Wesley, Reading, MA, USA, 2004. ISBN 0-13-124085-4. xxxii + 525 pp. LCCN ???? US$39.99, CAN$60.99.

Robbins:1999:UND

[44] Arnold Robbins. UNIX in a Nutshell: A Desktop Quick Reference for SVR4 and Solaris 7. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 1999. ISBN 1-56592-427-4. xvi + 598 pp. LCCN QA76.76.O63 R623 1999. US$24.95. URL http://www.oreilly.

com/catalog/unixnut3.

Robbins:2001:EAP

[45] Arnold Robbins. Effective AWK Programming. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2001. ISBN 0-596-00070-7. xxiv + 421 pp. LCCN QA76.73.A95 R63 2001. US$39.95.

URL http://www.oreilly.com/catalog/awkprog3/.

Robbins:2002:SAP

[46] Arnold Robbins. sed & awk Pocket Reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2002. ISBN 0-596-00352-8. 52 (est.) pp. LCCN ????

US$12.95. URL http://safari.oreilly.com/0596003528;   http://
www.oreilly.com/catalog/sedawkrepr2.

Robbins:2005:CSS

[47] Arnold Robbins and Nelson H. F. Beebe. Classic Shell Scripting. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00595-4. xxii + 534 pp. LCCN QA76.76.O63 R633 2005.

URL http://www.oreilly.com/catalog/shellsrptg/.

Rosenberg:1999:HKP

[48] Barry Rosenberg. Hands-On KornShell93 Programming. Addison-Wesley Longman, Harlow, Essex CM20 2JE, England, 1999. ISBN 0-201-31018-X. xxvi + 444 pp. LCCN QA76.73.K67 R67 1999.

Rosenblatt:2002:LKS

[49] Bill Rosenblatt and Arnold Robbins. Learning the Korn Shell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2002. ISBN 0-596-00195-9. xviii + 412 pp. LCCN QA76.73.K67 R68 2002. US$34.95.

Salus:1994:QCU

[50] Peter H. Salus. A quarter century of UNIX. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-54777-5. xii + 256 pp. LCCN QA76.76.O63 S342 1994.

Schneier:1996:ACP

[51] Bruce Schneier. Applied Cryptography: Protocols, Algorithms, and Source Code in C. Wiley, New York, NY, USA, second edition, 1996. ISBN 0-471-12845-7 (cloth), 0-471-11709-9 (paper). xxiii + 758 pp. LCCN QA76.9.A25 S35 1996.

Schneier:2000:SLD

[52] Bruce Schneier. Secrets and Lies: Digital Security in a Networked World. Wiley, New York, NY, USA, 2000. ISBN 0-471-25311-1. xv + 412 pp. LCCN QA76.9.A25 S352 2000. US$29.99.

Siever:2000:LND

[53] Ellen Siever et al., editors. Linux in a nutshell: a desktop quick reference. In a nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2000. ISBN 0-596-00025-1. xv + 797 pp. LCCN QA76.76.O63 L5459 2000; QA76.76.O63 L54595 2000.

Singh:1999:CBE

[54] Simon Singh. The code book: the evolution of secrecy from Mary, Queen of Scots, to quantum cryptography. Doubleday, New York, NY, USA, 1999. ISBN 0-385-49531-5. xiii + 402 pp. LCCN Z103 .S56 1999. US$24.95.

Stevens:1992:APU

[55] W. Richard Stevens. Advanced Programming in the UNIX Environment. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-56317-7. xviii + 744 pp. LCCN QA76.76.O63 S754 1992. US$52.25. The source codes and errata list are obtainable by anonymous ftp from ftp.uu.net (in /

published/books).

Sweet:2001:CCU

[56] Michael R. Sweet. CUPS: Common UNIX Printing System. SAMS Publishing, Indianapolis, IN, USA, 2001. ISBN 0-672-32196-3. xxvi + 622 pp. LCCN QA76.76.O63 S942 2001. US$44.99. URL http://www.cups.org;

 http://www.easysw.com/cups/pricing.html.

Unicode:2003:USV

[57] The Unicode Consortium. The Unicode Standard, Version 4.0. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-321-18578-1. xxxviii + 1462 pp.

LCCN QA268 .U545 2004. URL http://www.unicode.org/versions/

Unicode4.0.0/. Includes CD-ROM.

Vahalia:1996:UI

[58] Uresh Vahalia. UNIX Internals. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-101908-2. xxxiii + 601 pp. LCCN QA76.76.063V33 1996. US$47.75.

Walsh:1994:MTW

[59] Norman Walsh. Making TEX Work. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, April 1994. ISBN 1-56592-051-1. xxxvi + 483 pp. LCCN Z253.4.T47 W34 1994. US$29.95.

Watson:1990:FSC

[60] Thomas J. Watson Jr. Father Son & Co. — My Life at IBM and Beyond. Bantam Books, New York, NY, USA, 1990. ISBN 0-553-07011-8. xi + 468 pp. LCCN HD9696.C64 I4887 1990. Memoirs of IBM President Watson, the son of the founder of IBM.

Zimmermann:1995:OPU

[61] Philip R. Zimmermann. The Official PGP User’s Guide. MIT Press, Cambridge, MA, USA, 1995. ISBN 0-262-74017-6. xviii + 127 pp. LCCN TK5102.85 .Z56 1995. US$14.95. URL http://www-mitpress.mit.edu/

mitp/recent-books/comp/pgp-user.html.

Stevens:2005:APU

[62] W. Richard Stevens and Stephen A. Rago. Advanced programming in the Unix environment. Addison-Wesley, Reading, MA, USA, second edition, 2005. ISBN 0-201-43307-9 (hardcover). 960 (est.) pp. LCCN

QA76.76.O63 S754 2005. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/;   http://www.aw-bc.com/catalog/academic/product/
0,1144,0201433079,00.html;   http://www.loc.gov/catdir/toc/ecip059/
2005007943.html.